OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks. Number of pages: 50, 100, 150, 200 Cost : 1.55, 2.25, 2.95, 3.65 Which equation gives the cost, y, in terms of the number of pages, x? A. y = 0.014x + 49.15 B. y = 0.014x + 1.55 C. y = 0.014x + 0.85 D. y = 58.8x + 0.85

OpenStudy (ageta):

(3.65 -2.95 )/ (200-150) =m

OpenStudy (ageta):

soooooo like y-3.65= 0.014 (x-200) and you distribute 0.013x-2.8=y-3.65

OpenStudy (ageta):

add 3.65 on both sides

OpenStudy (ageta):

y=0.014x +0.85

OpenStudy (anonymous):

Thank you!

OpenStudy (ageta):

no problem